# Finite $N$ Hilbert Spaces of Bilocal Holography

### v1

@@ 1 Introduction
At finite $N$, one still has the question of non-perturbative completeness of the theory. For a theory with $K$ degrees of freedom with $K>N$ as a rule one has finite $N$ trace relations, whose implementation is of central importance. There were indications Jevicki:1991yi that the theory nevertheless applies generally, the fundamental reason being that the finite $N$ trace relations appear to be consistent with the collective representation, in particular they appear as null-relations preserved by the infinite space Hamiltonian. Recently deMelloKoch:2025ngs ; deMelloKoch:2025rkw ; deMelloKoch:2025eqt we have been conducting a systematic study of the space of invariants at finite $N$, and implementation of finite trace relations. The algebraic structure that emerges is succinctly given by the Hironaka decomposition, where the reduced set of invariants consists of invariants of the first kind (primary) which are freely acting and invariants of the second kind (secondaries) which have a constrained action. The fact that this algebraic structure can be implemented on the infinite collective space stems from the mathematical theory of Procesi Pr . The above algebraic reduction was seen to lead to reduction of the Hilbert space and it was most completely established on the example of theories enjoying an $S_{N}$ gauge invariance deMelloKoch:2025eqt where a $q$-deformed Hilbert space was established.
@@
The structure of Hilbert space of a theory is closely related to its finite temperature properties. In the case of vector/AdS/dS holography this is even more pertinent. The finite temperature CFT dual in AdS are black holes, and generally one has in matrix Sundborg:1999ue ; Aharony:2005bq ; Aharony:2003sx (and vector models Shenker:2011zf ; Amado:2016pgy ) a confinement/deconfinement phase transition. The importance of the relevant finite $N$ relations in this regard was first emphasized in Shenker:2011zf .
In the present work we present a comprehensive discussion of finite $N$ reduction of Hilbert space in bi-local collective theories. Full analysis of trace relations (following our matrix and $S_{N}$ studies) will show consistency with the overcomplete collective and geometric quantization scheme. This will be done both in bosonic and fermionic theories, with emphasis on the distinct structure of emergent Hilbert spaces in the two theories. We will then consider collective fields at finite temperature and show that the right state content is present. In particular a singlet space trace and partition function are evaluated.
@@ 3 Trace Relations and Reduction of Hilbert Space
respectively, for the bosonic and fermionic models. In comparison for the bosonic case Hilbert space we have a set of freely acting bosonic oscillators (giving the primary invariants in the Hilbert space, reflected in the denominator of (71)) and a finite (but exponentially large) set of secondary states reflected in the numerator of (71). The primaries generate a Fock space structure and lead to an infinite dimensional Hilbert space in bosonic theories.
For fermions, we only have secondary states, reflected in the fact that (72) has no denominator. Primary invariants (free oscillators) are absent, resulting in a finite Hilbert space for fermionic (and de Sitter) holography (based on $Sp(2N)$ fermions). In particular we emphasize that, although correlation functions of gauge-invariant operators in the bosonic $O(N)$ vector model are related to those of the fermionic $Sp(N)$ vector model by an analytic continuation $N\to-N$ Ramgoolam:1993hh ; Anninos:2011ui . Such a relation (and continuation) does not hold at the non-perturbative Hilbert space level  [^1] This implies that the proposal made in Anninos:2017eib is neither correct nor feasible.
@@ 4.2 Kähler Quantization
Bi-local fields become finite dimensional matrices of size $K$, belonging to a compact symmetric (Kahler) space
@@
This is manifold $M_{I}(K,K)$ in the classification of Berezin:1975 . Quantization on Kähler manifolds has been formulated by Berezin Berezin:1978sn . The Kähler scalar product defining the bi-local Hilbert space reads
@@ 5 Trace and Partition Function
$Z(\beta)$ $=$ $C(N,K){{\rm Vol}\,\Omega\over K!}\int\prod_{l=1}^{K}d\omega_{l}\,\,\Delta(\omega_{1}^{2},\omega_{2}^{2},\cdots,\omega_{K}^{2})^{2}\,\prod_{l}(1+\omega_{l}^{2})^{-2K-N}$ (141)
$\qquad\qquad\qquad\qquad\qquad\qquad\times\prod_{l}(1+x\omega^{2})^{N}$
@@
$C(N,K)^{-1}$ $=$ ${{\rm Vol}\,\Omega\over K!}\int\Delta(\omega_{1}^{2},\omega_{2}^{2},\cdots,\omega_{K}^{2})^{2}\prod_{l}(1+\omega^{2})^{-2N-K}\,$, (142)
@@
$Z(\beta)$ $=$ $\frac{{\rm Vol}\,\Omega}{K!}C(N,K)\int_{0}^{1}\prod_{l=1}^{K}dy_{i}\Delta(y_{1},y_{2},\cdots,y_{K})^{2}\prod_{i=1}^{K}\frac{(1-y_{i})^{N}}{(1-xy_{i})^{N}}$ (144)
@@
$I_{K,N}(x)\;=\;\int_{0}^{1}\cdots\int_{0}^{1}\Bigl(\Delta(y_{1},\dots,y_{K})\Bigr)^{2}\prod_{i=1}^{K}\frac{(1-y_{i})^{N}}{(1-xy_{i})^{N}}\,dy_{i}.$ (145)
@@
$I_{K,N}(x)=K!\,\det\!\Bigl[\mu_{i+j}\Bigr]_{i,j=0}^{K-1},\qquad\mu_{m}=\int_{0}^{1}y^{m}\,w(y)\,dy=\int_{0}^{1}y^{m}\frac{(1-y)^{N}}{(1-xy)^{N}}\,dy.$ (153)
@@
$_{2}F_{1}(A,B;C;x)=\frac{\Gamma(C)}{\Gamma(B)\Gamma(C-B)}\int_{0}^{1}t^{B-1}(1-t)^{C-B-1}(1-xt)^{-A}dt.$ (155)
@@
$A=N,\qquad B=m+1,\qquad C=m+N+2.$ (156)
Then CB1=NC-B-1=N, and the integral representation becomes exactly
$\int_{0}^{1}t^{m}(1-t)^{N}(1-xt)^{-N}dt=\frac{\Gamma(m+1)\Gamma(N+1)}{\Gamma(m+N+2)}\,{}_{2}F_{1}\!\bigl(N,m+1;m+N+2;x\bigr)$ (157)
@@
$Z(\beta)=N_{Z}\,\det\!\big[B(i+j+1,N+1)\,{}_{2}F_{1}\bigl(N,i+j+1;N+i+j+2;x\bigr)\big]_{i,j=0}^{K-1}.$ (159)
@@ 6 Conclusions
We have discussed in this work two descriptions of the Hilbert space for finite $N$ singlets in vector theories: (i) the bilocal formulation, in which the space of configurations is viewed as a Kähler phase space and quantized using geometric quantization, and (ii) the direct construction in terms of gauge-invariant operators, whose counting is captured by the Molien–Weyl formula. We showed that these perspectives are equivalent, in the sense that the number of independent invariants coincides with the dimension of the Hilbert space, obtained through (overcomplete) geometric quantization. Finally a third specification of the Hilbert space is introduced, based on group of invariants, the details of this will be presented in a subsequent publication.
@@ Acknowledgements
The work of RdMK was supported by a start up research fund of Huzhou University, a Zhejiang Province talent award and by a Changjiang Scholar award. J.Y. was supported by the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) (RS-2022-NR069038) and by the Brain Pool program funded by the Ministry of Science and ICT through the National Research Foundation of Korea (RS-2023-00261799).

### v2

@@ 1 Introduction
At finite $N$, one still has the question of non-perturbative completeness of the theory. For a theory with $K$ degrees of freedom with $K>N$ as a rule one has finite $N$ trace relations, whose implementation is of central importance. There were indications Jevicki:1991yi that the theory nevertheless applies generally, the fundamental reason being that the finite $N$ trace relations appear to be consistent with the collective representation, in particular they appear as null-relations preserved by the infinite space Hamiltonian. Recently deMelloKoch:2025ngs ; deMelloKoch:2025rkw ; deMelloKoch:2025eqt we have been conducting a systematic study of the space of invariants at finite $N$, and implementation of finite trace relations. The algebraic structure that emerges is succinctly given by the Hironaka decomposition, where the reduced set of invariants consists of invariants of the first kind (primary) which are freely acting and secondary invariants ,whose number is shown to grow exponentially (in powers of N). These secondary states, could be termed “miraculous” as they surpass the finite $N$ cut off bounds. The fact that this algebraic structure can be implemented on the infinite collective space stems from the mathematical theory of Procesi Pr . The above algebraic reduction was seen to lead to reduction of the Hilbert space and it was most completely established on the example of theories enjoying an $S_{N}$ gauge invariance deMelloKoch:2025eqt where a $q$-deformed Hilbert space was established.
@@
The structure of Hilbert space of a theory is closely related to its finite temperature properties. In the case of vector/AdS/dS holography this is even more pertinent. The finite temperature CFT with dual AdS black holes, one generally sees a confinement /deconfinement phase transition appearing both in matrix Sundborg:1999ue ; Aharony:2005bq ; Aharony:2003sx and vector models Shenker:2011zf ; Amado:2016pgy ), where the importance of finite $N$ relations for the transition is first emphasized in Shenker:2011zf .
Finite-NN effects also play a central role in the description of black hole microstates, particularly in the discovery of the fortuity mechanism Chang:2022mjp ; Choi:2022caq ; Choi:2023znd ; Chang:2023zqk ; Choi:2023vdm ; Chang:2024zqi ; Kim:2025vup : the microstates of $\frac{1}{16}$-BPS black holes fall into two classes, monotone and fortuitous. Monotone states remain BPS at all $N$, while fortuitous states lose their BPS nature above a critical $N$. In this there is definitely a similarity with the states that we have established, and hopefully future work will illuminate the connection.
In the present work we present a comprehensive discussion of finite $N$ reduction of Hilbert space in bi-local collective theories. Full analysis of trace relations (following our matrix and $S_{N}$ studies) will show consistency with the overcomplete collective and geometric quantization schemeDas:2012dt for defining the Hilbert space of de Sitter/Sp($2N$) holography. We will perform a systematic discussion both in bosonic and fermionic theories, with emphasis on the distinct structure of emergent Hilbert spaces in the two cases. We will then consider collective fields at finite temperature and show that the right state content is present. In particular a singlet space trace and partition function are evaluated.
@@ 3 Trace Relations and Reduction of Hilbert Space
respectively, for the bosonic and fermionic models. In comparison for the bosonic case Hilbert space we have a set of freely acting bosonic oscillators (giving the primary invariants in the Hilbert space, reflected in the denominator of (71)) and a finite (but exponentially large) set of secondary states reflected in the numerator of (71). The appearance of these states is definitelly ”miraculous” as they surpass the finite N cutoff bounds . Our analysis has established their exponential (in powers of N) growth .As such they play a central role in thermodynamics and are responsible for finite Entropy.
For fermions, we only have secondary states, reflected in the fact that (72) has no denominator. Primary invariants (free oscillators) are absent, resulting in a finite Hilbert space for fermionic (and de Sitter) holography (based on $Sp(2N)$ fermions). In particular we emphasize that, although correlation functions of gauge-invariant operators in the bosonic $O(N)$ vector model are related to those of the fermionic $Sp(N)$ vector model by an analytic continuation $N\to-N$ Ramgoolam:1993hh ; Anninos:2011ui . Such a relation (and continuation) does not hold at the non-perturbative Hilbert space level.
@@ 4.2 Kähler Quantization
Bi-local fields become finite dimensional matrices of size $K$, and due to the quadratic (matrix) constraints form a compact symmetric (Kahler) space
@@
For the hermitean matrix case we have the manifold $M_{I}(K,K)$ in the classification of Berezin:1975 with the Kähler scalar product given by
@@ 5 Trace and Partition Function
$Z(\beta)$ $=$ $C(N,K){{\rm Vol}\,\Omega\over K!}\int\prod_{l=1}^{K}d\omega_{l}\,\,\Delta^{2}(\omega_{1}^{2},\omega_{2}^{2},\cdots,\omega_{K}^{2})\,\prod_{l}(1+\omega_{l}^{2})^{-2K-N}$ (141)
$\qquad\qquad\qquad\qquad\qquad\qquad\times\prod_{l}(1+x\omega_{l}^{2})^{N}$
@@
$C(N,K)^{-1}$ $=$ ${{\rm Vol}\,\Omega\over K!}\int\Delta^{2}(\omega_{1}^{2},\omega_{2}^{2},\cdots,\omega_{K}^{2})\prod_{l}(1+\omega_{l}^{2})^{-2K-N}\,$, (142)
@@
$Z(\beta)$ $=$ $\frac{{\rm Vol}\,\Omega}{2^{K}K!}C(N,K)\int_{0}^{1}\prod_{l=1}^{K}dy_{i}\Delta^{2}(y_{1},y_{2},\cdots,y_{K})\prod_{i=1}^{K}(1-(1-x)y_{i})^{N}$ (144)
@@
$I_{K,N}(x)\;=\;\int_{0}^{1}\cdots\int_{0}^{1}\Bigl(\Delta(y_{1},\dots,y_{K})\Bigr)^{2}\prod_{i=1}^{K}(1-(1-x)y_{i})^{N}\,dy_{i}.$ (145)
@@
$I_{K,N}(x)=K!\,\det\!\Bigl[\mu_{i+j}\Bigr]_{i,j=0}^{K-1},\qquad\mu_{m}=\int_{0}^{1}y^{m}\,w(y)\,dy=\int_{0}^{1}y^{m}(1-(1-x)y_{i})^{N}\,dy.$ (153)
@@
$_{2}F_{1}(A,B;C;\xi)=\frac{\Gamma(C)}{\Gamma(B)\Gamma(C-B)}\int_{0}^{1}t^{B-1}(1-t)^{C-B-1}(1-\xi t)^{-A}dt.$ (155)
@@
$A=-N,\qquad B=m+1,\qquad C=m+2.$ (156)
Then the integral representation becomes
$\int_{0}^{1}t^{m}(1-(1-x)t)^{-N}dt=\frac{1}{m+1}\,{}_{2}F_{1}\!\bigl(-N,m+1;m+2;1-x\bigr)$ (157)
@@
$Z(\beta)=N_{Z}\,\det\Big[\frac{1}{i+j+1}\,{}_{2}F_{1}\bigl(-N,i+j+1;i+j+2;1-x\bigr)\Big]_{i,j=0}^{K-1}.$ (159)
@@ 6 Conclusions
We have discussed in this work two descriptions of the Hilbert space for finite $N$ singlets in vector theories: (i) the bilocal formulation, in which the space of configurations is viewed as a Kähler phase space and quantized using geometric quantization, and (ii) the direct construction in terms of gauge-invariant operators, whose counting is captured by the Molien–Weyl formula. We showed that these perspectives are equivalent, in the sense that the number of independent invariants coincides with the dimension of the Hilbert space [^1] The proposal for the Hilbert space made in Anninos:2017eib does not agree with the finite $N$ Hilbert space of either bosons or fermions. In particular, that proposal misses the secondary class of states., obtained through (overcomplete) geometricDas:2012dt quantization. Finally a third specification of the Hilbert space is introduced, based on group of invariants; the details of this will be presented in the companion publication Us2 .
@@ Acknowledgements
We would like to acknowledge discussions with Sumit Das and Junjie Zheng. We are grateful to Junjie Zheng for correcting an error in the earlier version. The work of RdMK was supported by a start up research fund of Huzhou University, a Zhejiang Province talent award and by a Changjiang Scholar award. J.Y. was supported by the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) (RS-2022-NR069038) and by the Brain Pool program funded by the Ministry of Science and ICT through the National Research Foundation of Korea (RS-2023-00261799).